a.      Cisplatin or Platinol has a chemical PtCl2(NH3)2 that is used in the treatment of certain cancers. Calculate (a) the amount (mol) of compound in 285.3 g of cisplatin; (b) the number of hydrogen atoms in 0.98 mol of cisplatin.



1
Expert's answer
2022-03-17T15:02:43-0400

Solution:

(a):

The molar mass of cisplatin (PtCl2(NH3)2) is 300.051 g/mol

Mass of PtCl2(NH3)2 is 285.3 g

Therefore,

Moles of cisplatin = (285.3 g) × (1 mol / 300.051 g) = 0.9508 mol

The moles of cisplatin is 0.9508 mol


(b):

PtCl2(NH3)2 ⟶ 6H

According to stoichiometry:

1 mol of cisplatin contains 6 mol of H atoms

Thus, 0.98 mol of cisplatin contains (0.98×6) = 5.88 mol of H atoms

One mole of any substance contains 6.022×1023 particles

Therefore,

Number of H atoms = (5.88 mol) × (6.022×1023 atoms / 1 mol) = 3.54×1024 atoms

3.54×1024 hydrogen atoms in 0.98 mol of cisplatin